how much water should be added to 60 litres of milk containing 20% cream in order to get milk containing 15% cream

Found 2 solutions by TimothyLamb, mananth:
Answer by TimothyLamb(4379)

60(0.20) + V(0.00) = (V + 60)(0.15)
12 = (0.15)V + 9
(0.15)V = 3
V = 3/(0.15)
---
V = 20 liters
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php

Answer by mananth(16946)

--- percent ---------------- Amount
milk 20 ---------------- 60 liters
water 0 ---------------- x liters
Mixture 15 ---------------- 60 + x liters

20 * 60 + 0 x = 15 ( 60 + x )
1200 + 0 x = 900 + 15 x
0 x -15 x = -1200 + 900
-15 x = -300
/ -15
x = 20 liters water

m.ananth@hotmail.ca